Calculate the concentration of each standard in terms of ppm iron. FW= 55.845 g/mol. Please show your work. First prepare standard solution from a standard Fe stock of 0.13 M. Make 100x dilution (1 mL of stock into 100 mL of water) Then, using a pipet deliver the following volumes of your Fe standard diluted solution into 10 mL volumetric flasks: 0 microliters, 150 micro liters, 300 microliters, 450 microliters, and 600 microliters.
Answer #1

Concentration of each standard in terms of ppm of iron is calculated as follows:
Answer #2

Given:

  • Stock Fe solution concentration: 0.13 M

  • Dilution: 1 mL stock → 100 mL (100x dilution)

  • Final volumes: 10 mL (in volumetric flasks)

  • Volumes of diluted standard added: 0 µL, 150 µL, 300 µL, 450 µL, 600 µL

  • Molar mass of Fe (FW): 55.845 g/mol



Step 1: Prepare the Diluted Standard Solution

  1. Diluted Fe concentration (100x dilution):

    Mdiluted=0.13M100=0.0013M

  2. Convert to ppm (mg/L):

    ppm=M×FW×1000=0.0013×55.845×1000=72.60ppm


Step 2: Calculate Fe Concentration in Each 10 mL Flask

General Formula:

Cfinal=(Vadded(µL)10,000µL)×72.60ppm

(Since 10 mL = 10,000 µL)

Volume Added (µL)CalculationFinal Concentration (ppm Fe)
000.00
150(150/10,000)×72.601.09
300(300/10,000)×72.602.18
450(450/10,000)×72.603.27
600(600/10,000)×72.604.36


Answer:

StandardFe Concentration (ppm)
Blank0.00
150 µL1.09
300 µL2.18
450 µL3.27
600 µL4.36


answered by: anonymous